OpenStudy (thewalker1132):

solve the system by substitution 2x-y+z=-4 z=5 -2x+3y-z=-10 a)(-8,7,5) b)(-8,-7,5) c)(8,-7,5) d)(-8,-7,-5)

OpenStudy (solomonzelman):

you are given that z=5 from the beginning , so you can go ahead and substitute this into the remaining two equations.

OpenStudy (solomonzelman):

2x-y+5=-4 -2x+3y-5=-10 2x - y = -9 -2x + 3y = -5

OpenStudy (solomonzelman):

then add the two equations and that will help you solve for y.